A team of 2 men and 5 women completed one –fourth of a job in 3 days. After 3 days another man joined the team and they took 2 days to complete another one –fourth of the job. How many men can complete the whole job in 4 days?
Answer by stanbon(75887)

A team of 2 men and 5 women completed one –fourth of a job in 3 days. After 3 days another man joined the team and they took 2 days to complete another one –fourth of the job. How many men can complete the whole job in 4 days?
----
Let:
men rate = m
woman rate = w
----
2m + 5w = 1/12
3(2m+5w) + 2(3m+5w) = 1/2
----
3(1/12) + 2(m+(1/12)) = 1/2
1/4 + 2m + (1/6) = 1/2
2m + (5/12) = 6/12
2m = 1/12
m = 1/24
----
One man can do the job in 24 days.
6 men can do the job in 4 days.
----
cheers,
Stan H.
----------------